# Medical Spoken Named Entity Recognition

### v2

@@ Abstract
Spoken Named Entity Recognition (NER) aims to extracting named entities from speech and categorizing them into types like person, location, organization, etc. In this work, we present VietMed-NER - the first spoken NER dataset in the medical domain. To our best knowledge, our real-world dataset is the largest spoken NER dataset in the world in terms of the number of entity types, featuring 18 distinct types. Secondly, we present baseline results using various state-of-the-art pre-trained models: encoder-only and sequence-to-sequence. We found that pre-trained multilingual models XLM-R outperformed all monolingual models on both reference text and ASR output. Also in general, encoders perform better than sequence-to-sequence models for the NER task. By simply translating, the transcript is applicable not just to Vietnamese but to other languages as well. All code, data and models are made publicly available here.
@@
Khai Le-Duc∗1,3,4,5, David Thulke1,2, Hung-Phong Tran7, Long Vo-Dang8, Khai-Nguyen Nguyen9, Truong-Son Hy5,6, Ralf Schlüter1,2 1Lehrstuhl Informatik 6 - Machine Learning and Human Language Technology Group, Computer Science Department, RWTH Aachen University, 52074 Aachen, Germany 2AppTek GmbH, 52062 Aachen, Germany 3University of Toronto, Canada 4University Health Network, Canada 5FPT Software AI Center, Vietnam 6Indiana State University, United States 7Hanoi University of Science and Technology, Vietnam 8University of Cincinnati, United States 9College of William and Mary, United States duckhai.le@mail.utoronto.ca TruongSon.Hy@indstate.edu, {thulke,schlueter}@hltpr.rwth-aachen.de
(*)(*)footnotetext: Work partially done during the bachelor thesis at Lehrstuhl Informatik 6 - Machine Learning and Human Language Technology Group, Computer Science Department, RWTH Aachen University
@@ 1 Introduction
Named Entity Recognition (NER) targets extracting named entities (NE) from text and categorizing them into types like person, location, organization, etc. Initially studied in written language, recent attention has turned to studying spoken NER Cohn et al. (2019); Shon et al. (2022), which aims to extract semantic information from speech. However, spoken NER has limited literature compared to NER on written text data Yadav et al. (2020).
Spoken NER is particularly challenging, firstly due to the impact of word segmentation on results. The Vietnamese vocabulary poses difficulties with numerous confused monosyllabic and polysyllabic words. For instance, the word "đường" alone could denote "sugar" (chemical), "street" (location), or be part of a compound word like "đường tiêu hóa" - "gastrointestinal" (anatomy). This confusion has also been reported in Chinese spoken NER by Chen et al. (2022). Secondly, data quality control and annotation consistency have been problematic, with some entities tagged in one sentence but not in others, and full NEs inconsistently tagged as multiple sub-NEs Huyen and Luong (2016); Nguyen et al. (2018, 2020); Truong et al. (2021). Thirdly, obtaining accurate medical NER from natural speech is challenging due to the lack of punctuation, speech disfluency, context, and the complexity of medical terms.
As for the medical domain, to our best knowledge, there is no dataset available for medical spoken NER. The only related work we found, Cohn et al. (2019), published a NER evaluation benchmark using an English general-domain conversational dataset, Switchboard Godfrey et al. (1992) and Fisher Cieri et al. (2004), for the task of audio de-identification specifically targeting Personal Health Identifiers.
To address this gap, we introduce VietMed-NER, a medical spoken NER dataset built on the real-world medical Automatic Speech Recognition (ASR) dataset VietMed Le-Duc (2024), featuring 18 medically-defined entity types. The dataset has the potential to serve as a foundational task in various real-world applications, such as: error correction for medical ASR output Mani et al. (2020), audio de-identification. By simply translating, the transcript is applicable not just to Vietnamese but to other languages as well, such as: privacy-preserved machine learning (e.g. for model training and public dataset creation), and virtual assistant (e.g. search engine Rüd et al. (2011), classifying content for news providers Kumaran and Allan (2004), and recommending content Koperski et al. (2017)). Our contributions are as follows:
@@
  • We present our annotation scheme, called Recursive Greedy Mapping, intending to improve annotation speed and quality in the medical domain.
  • We present baselines on several state-of-the-art pre-trained models and analyze the results.
All code, data (English-translated and Vietnamese) and models are published online [^1] https://github.com/leduckhai/MultiMed.
@@ 2.1 Data Collection
Entity Type and Definition Train Dev Test All
Total Uni. Total Uni. Total Uni. Total Uni.
AGE 447 43 108 25 611 83 1166 151
Age of a person
GENDER 202 30 46 15 451 33 699 78
Gender of a person
OCCUPATION 543 32 133 16 562 43 1238 91
Job of a person
LOCATION 284 66 76 31 317 75 677 172
Locations and places
ORGANIZATION 19 14 2 2 58 23 79 39
Organizations
DISEASESYMTOM 2699 518 683 209 1334 357 4716 1084
Diseases and symptoms
DRUGCHEMICAL 1054 255 263 104 684 136 2001 495
Bio-chemical substances and drugs
FOODDRINK 243 77 48 26 247 43 538 146
Food and beverage
ORGAN 1827 252 444 122 1190 172 3461 546
Anatomical features, e.g. organs, cells
PERSONALCARE 353 114 82 38 95 10 530 162
Personal care, e.g. hygiene routines, skin care
DIAGNOSTICS 371 53 91 25 292 36 754 114
Diagnostic procedures, e.g. lab tests, imaging
TREATMENT 726 69 174 25 230 17 1130 111
Non-surgical treatment, e.g. rehab., injection
SURGERY 197 29 55 13 270 37 522 79
Surgical procedures, e.g. implants, neurosurgery
PREVENTIVEMED 341 53 80 25 18 6 439 84
Preventive medicine
MEDDEVICETECHNIQUE 324 84 67 30 603 144 994 258
Medical devices, instruments, and techniques
UNITCALIBRATOR 800 155 215 75 251 106 1266 336
Medical calibration, e.g. number of doses, calories
TRANSPORTATION 5 2 3 3 27 10 35 15
Means of transportation
DATETIME 674 155 159 65 657 133 1490 353
Date and time
#Entities in total 11109 2001 2729 849 7897 1464 21735 4314
#Sentences 4620 1150 3500 9270
Table 1: Entity definition and its statistics in our dataset. "Uni." means the number of unique entities, which equals to the length of gazetteer list.
We chose the VietMed dataset Le-Duc (2024), the world’s largest and most generalizable and publicly-available medical ASR dataset, for annotating NEs.
@@ 2.2 Annotation Process
Annotation of medical NEs from real-world speech is challenging because of the missing punctuation, special characters and capitalized words in ASR transcript, disfluencies and required medical knowledge. Entirely manual annotation of NEs like in VLSP dataset Huyen and Luong (2016); Nguyen et al. (2018, 2020) and PhoNER_COVID19 Truong et al. (2021) requires a large number of working hours, not to mention the difficulties in quality control and inconsistency as we found in their corpora. This inconsistency includes: i) Some entities tagged in a sentence are not also tagged in other sentences, and ii) Full NEs are inconsistently tagged as multiple sub-NEs.
To tackle these problems, we present a human-machine annotation approach, called Recursive Greedy Mapping. The method is described below:
@@
    Annotate and categorize some initial entities, then add them to a gazetteer list.
@@
    Sort entities by character length from highest to lowest, to distinguish between sub-NEs and full NEs, ensuring full NEs are mapped before sub-NEs. Time complexity = $O(k\cdot\log(k))$ where $k$ is the number of NEs.
@@
In short, Recursive Greedy Mapping speeds up annotation by allowing annotators to foresee possible NEs in upcoming utterances based on previously annotated ones. Annotators can accept or reject these suggestions, saving time with correct suggestions and easily ignoring incorrect ones. Unlike training a seed model with a gazetteer, which requires initial training time and may be unreliable for low-to-mid resource languages, our method avoids these issues and eliminates the need to correct incorrect NEs.
@@ 2.3 Data Quality Control
We created initial annotation guidelines (see Appendix) and began annotating the corpus. Two developers, one with a medical background, independently annotated the corpus. Then, we held a discussion session to resolve conflicts, address complex cases, and refine the guidelines. Two other developers perform quality control using the guidelines and the annotated corpus. We consistently revisited each sentence in the entire corpus multiple times.
@@ 2.4 Data Statistics
Most NER datasets have a very small number of entities in test sets compared to train and dev set Huyen and Luong (2016); Truong et al. (2021); Chen et al. (2022). But we want to leverage the capabilities of large pre-trained models which were trained on vast amounts of unlabeled text data, resulting in good representations. Therefore, we leave the test set to be larger than train and dev sets. However, this split caused slight data imbalance.
@@ 3 Experimental Setups
We employ the 2-stage pipeline for spoken NER: A hybrid ASR model transcribes audio into text and then the transcribed text is fed into a NER model.
3.1 F1 Score
We employed seqeval [^2] https://github.com/chakki-works/seqeval framework commonly used as a default evaluation framework by HuggingFace. However, this framework only works for NER on reference text. Therefore, we also employed the F1 score calculation by Shon et al. (2023) by using the SLUE toolkit [^3] https://github.com/asappresearch/slue-toolkit. This F1 score evaluates an unordered list of named entity phrase and tag pairs predicted for each sentence. Our proposed modification of SLUE was also used and presented in the Appendix.
@@ 3.2 ASR Models
We employed 2 best baseline models fine-tuned for ASR task on VietMed published by Le-Duc (2024): monolingual w2v2-Viet and cross-lingual XLSR-53-Viet. w2v2-Viet model was pre-trained on 1204h of Vietnamese data. XLSR-53-Viet model was pre-trained on 1204h of Vietnamese with XLSR-53 Conneau et al. (2021) as initialization. Both of them have the same number of parameters (118M) and were fine-tuned on the same train set. Their WERs on test set were 29.0% and 28.8% respectively.
@@ 3.3 NER Models
Table 2 shows the statistics of various pre-trained monolingual and multilingual models to fine-tune on our dataset. Monolingual: PhoBERT_base, PhoBERT_large, PhoBERT_base-v2 Nguyen and Nguyen (2020); ViDeBERTa_base Tran et al. (2023), BARTpho Tran et al. (2021), ViT5 Phan et al. (2022), Multilingual: XML-R_base, XML-R_large Conneau et al. (2020). To our best knowledge, these are the best pre-trained models that achieved state-of-the-art results on various downstream tasks in the Vietnamese language, including NER.
We used HuggingFace Wolf et al. (2019) for fine-tuning pre-trained models on the NER task. All our NER experiments were done using the default hyperparameters from HuggingFace. Details of the hyperparameters are shown in the Appendix.
@@ 4 Experimental Results
Table 3: NER results (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro F1 score. Results by entity types are shown in Appendix.
NER ASR Prec Rec. F1
@@
Table 4: NER results (in percent) on ASR output of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro F1 score. Results by entity types are shown in Appendix.
Table 3 and 4 show results of NER using various pre-trained models. The pre-trained monolingual model PhoBERT_base-v2 outperformed other monolingual models, at 61.99% of F1 score on reference text, and 51.29% on ASR output. Even though it has fewer parameters than PhoBERT_large, it achieved equal performance probably due to the larger amount of pre-training data. All pre-trained multilingual models outperformed monolingual models. This gap can be explained by the larger amount of pre-training data (2.5TB of multilingual data for XLM-R compared to only 140GB of monolingual data for PhoBERT_base-v2). Furthermore, encoder-based models perform, on average, better than seq2seq models, which could be due to the nature of seq2seq being generative, thus less suitable for NER. From the results from Table 3 and 4, we observe that there is a performance drop in all models when evaluated on ASR transcripts. This behavior is expected due to the noisy nature of ASR data.

### v3

@@ Abstract
Spoken Named Entity Recognition (NER) aims to extract named entities from speech and categorise them into types like person, location, organization, etc. In this work, we present VietMed-NER - the first spoken NER dataset in the medical domain. To our knowledge, our Vietnamese real-world dataset is the largest spoken NER dataset in the world regarding the number of entity types, featuring 18 distinct types. Furthermore, we present baseline results using various state-of-the-art pre-trained models: encoder-only and sequence-to-sequence; and conduct quantitative and qualitative error analysis. We found that pre-trained multilingual models generally outperform monolingual models on reference text and ASR output and encoders outperform sequence-to-sequence models in NER tasks. By translating the transcripts, the dataset can also be utilised for text NER in the medical domain in other languages than Vietnamese. All code, data and models are publicly available [^1] https://github.com/leduckhai/MultiMed/tree/master/VietMed-NER.
@@
Khai Le-Duc1,2, David Thulke4,5, Hung-Phong Tran3, Long Vo-Dang7, Khai-Nguyen Nguyen8, Truong-Son Hy6, Ralf Schlüter4,5 1University of Toronto, Canada 2University Health Network, Canada 3Hanoi University of Science and Technology, Vietnam 4Machine Learning and Human Language Technology Group, RWTH Aachen University, Germany 5AppTek GmbH, Germany 6University of Alabama at Birmingham, United States 7University of Cincinnati, United States 8College of William and Mary, United States duckhai.le@mail.utoronto.ca thy@uab.edu, {thulke,schlueter}@hltpr.rwth-aachen.de
@@ 1 Introduction
Named Entity Recognition (NER) targets extracting named entities (NE) from text and categorizing them into types like person, location, organization, etc. Initially studied in written language, recent attention has turned to study spoken NER Cohn et al. (2019); Shon et al. (2022), which aims to extract semantic information from speech. However, spoken NER has limited literature compared to NER on written text data Yadav et al. (2020).
Spoken NER is particularly challenging, firstly due to the impact of word segmentation on results. The medical vocabulary poses difficulties with numerous confused monosyllabic and polysyllabic words. For instance, the word "đường" alone could denote "sugar" (chemical), "street" (location), or be part of a compound word like "đường tiêu hóa" - "gastrointestinal" (anatomy). This confusion has also been reported in Chinese spoken NER by Chen et al. (2022). Further, data quality control and annotation consistency have been problematic, with some entities tagged in one sentence but not in others, and full NEs inconsistently tagged as multiple sub-NEs Huyen and Luong (2016); Nguyen et al. (2018, 2020); Truong et al. (2021). Finally, obtaining accurate medical NER from natural speech is challenging due to the lack of punctuation Ertopçu et al. (2017), speech disfluencies Kim and Woodland (2000), context, and the complexity of medical terms.
As for the medical domain, to the best of our knowledge, there is no dataset available for medical spoken NER. The only related work we found, Cohn et al. (2019), published a NER evaluation benchmark using an English general-domain conversational dataset, Switchboard Godfrey et al. (1992) and Fisher Cieri et al. (2004), for the task of audio de-identification specifically targeting Personal Health Identifiers.
To address this gap, we introduce VietMed-NER, a medical spoken NER dataset built on the real-world medical Automatic Speech Recognition (ASR) dataset VietMed Le-Duc (2024), featuring 18 medically-defined entity types. In the era of the advanced in-context learning capabilities of Large Language Models (LLMs) and human-level text-to-speech technologies, the dataset, with entity positional labels maintained during translation, is applicable not only to Vietnamese but also to other languages (see Appendix C). This enables various real-world applications, including: search engines Rüd et al. (2011), content classification for news providers Kumaran and Allan (2004), medical ASR error correction Mani et al. (2020), audio de-identification Cohn et al. (2019) and content recommendation systems Koperski et al. (2017).
Our contributions are as follows:
@@
  • We present baselines on several state-of-the-art pre-trained models
  • We conduct quantitative and qualitative error analysis for medical spoken NER in Vietnamese
All code, data and models are published online11footnotemark: 1.
Definition Train Dev Test All
Total Uni. Total Uni. Total Uni. Total Uni.
AGE Age of a person 447 43 108 25 611 83 1166 151
GENDER Gender of a person 202 30 46 15 451 33 699 78
JOB Job of a person 543 32 133 16 562 43 1238 91
LOCATION Locations and places 284 66 76 31 317 75 677 172
ORGANIZATION Organizations 19 14 2 2 58 23 79 39
DISEASESYMTOM Symptoms and diseases 2699 518 683 209 1334 357 4716 1084
DRUGCHEMICAL Bio-chemical substances and drugs 1054 255 263 104 684 136 2001 495
FOODDRINK Food and beverage 243 77 48 26 247 43 538 146
ORGAN Anatomical features, e.g. organs, cells 1827 252 444 122 1190 172 3461 546
PERSONALCARE Personal care, e.g. hygiene routines, skin care 353 114 82 38 95 10 530 162
DIAGNOSTICS Diagnostic procedures, e.g. lab tests, imaging 371 53 91 25 292 36 754 114
TREATMENT Non-surgical treatment, e.g. rehab., injection 726 69 174 25 230 17 1130 111
SURGERY Surgical procedures, e.g. implants, neurosurgery 197 29 55 13 270 37 522 79
PREVENTIVEMED Preventive medicine 341 53 80 25 18 6 439 84
MEDDEVICETECHNIQUE Medical devices, instruments, and techniques 324 84 67 30 603 144 994 258
UNITCALIBRATOR Medical calibration, e.g. number of doses, calories 800 155 215 75 251 106 1266 336
TRANSPORTATION Means of transportation 5 2 3 3 27 10 35 15
DATETIME Date and time 674 155 159 65 657 133 1490 353
#Entities in total 11109 2001 2729 849 7897 1464 21735 4314
#Sentences 4620 1150 3500 9270
Table 1: Entity definition and its statistics in our dataset. "Uni." means the number of unique entities.
2 Related Works
Traditionally, spoken NER has been done using a pipeline methodology, also known as cascaded approach, starting with an ASR stage, followed by NER applied to the generated transcriptions Jannet et al. (2017); Benaicha et al. (2024). Another variant of the cascaded approach involves embedding specific entity expressions into the lexicon, thereby improving the language model’s ability to accurately recognize these expressions Hatmi et al. (2013).
Besides, end-to-end NER has recently garnered some attention within the research community. This approach seeks to optimize ASR and NER processes simultaneously, offering a potentially more efficient alternative to traditional pipeline methods by harnessing the ability of trainable acoustic features. However, its accuracy advantage over the cascaded approach remains a subject of debate, and the end-to-end training setup introduces additional complexity Tomashenko et al. (2019); Yadav et al. (2020).
@@ 3.1 Data Collection
We chose the VietMed dataset Le-Duc (2024), the world’s largest publicly available medical ASR dataset, for annotating NEs.
The original dataset is in Vietnamese. We annotate the Vietnamese version with the methodology described in Section 3.2 and automatically translate the transcripts to English together with transferring the NE annotation.
@@ 3.2 Annotation Process
The annotation of medical NEs from real-world speech is challenging because of the missing punctuation, special characters and capitalized words in ASR transcripts, disfluencies and required medical knowledge. Entirely manual annotation of NEs like in VLSP dataset Huyen and Luong (2016); Nguyen et al. (2018, 2020) and PhoNER_COVID19 Truong et al. (2021) requires a large number of working hours, not to mention the difficulties in quality control and inconsistency as we found in their corpora. These inconsistencies include: i) Some entities tagged in one sentence are not tagged in another sentence, and ii) Full NEs are inconsistently tagged as multiple sub-NEs. The best approach to tag nested NEs is the subject of ongoing debate Muis and Lu (2017); Li et al. (2021a). For simplicity, higher consistency and to reduce the annotation effort, we only annotate the largest and outermost full entity span.
Moreover, using fine-tuned models for pre-tagging doesn’t apply to specific medical entity types. Similarly, using prompt engineering with large language models like GPT-4 for pre-tagging did not achieve acceptable accuracy. Training a seed model with a gazetteer list requires initial training time, subsequent repetitive training schedules, and may prove unreliable due to its statistical reliance on a small amount of low-resource data Kozareva (2006).
To tackle these problems, we conduct a human-machine annotation approach, as described below:
@@
    Annotate and categorize a set of initial entities, then add them to a gazetteer list.
@@
    Sort entities by character length from highest to lowest, to distinguish between sub-NEs and full NEs, ensuring full NEs are mapped before sub-NEs. Time complexity = $O(k\cdot\log(k))$ where $k$ is the number of NEs. For example, "tooth pain" should be mapped before "pain".
@@
We experience faster annotation by allowing annotators to foresee possible NEs in upcoming utterances based on previously annotated ones. Annotators can accept or reject these suggestions, saving time with correct suggestions and easily ignoring incorrect ones. Unlike training a seed model with a gazetteer, which requires initial training time and may be unreliable for low-to-mid resource languages, our method avoids these issues and eliminates the need to correct incorrect NEs.
@@ 3.3 Data Quality Control
We created initial annotation guidelines (see Appendix A) and began annotating the corpus. Two developers, one with a medical background, independently annotated the corpus. Then, we held a discussion session to resolve conflicts, address complex cases, and refine the guidelines. Two other developers perform quality control using the guidelines and the annotated corpus. We consistently revisited each sentence in the entire corpus multiple times. This data quality control process is inspired by Tran et al. (2022).
3.4 Data Splitting
Most NER datasets have a very small number of entities in their test sets compared to train and dev set Huyen and Luong (2016); Truong et al. (2021); Chen et al. (2022). However, we want to leverage the capabilities of large pre-trained models which are trained on vast amounts of unlabeled text data, resulting in good representations. Therefore, we focus on creating a large test set to obtain more statistically significant evaluation results and keep the training set relatively small in comparison.
@@ 4 Experimental Setups
We employ the cascaded (two-stage) pipeline for spoken NER: A hybrid ASR model transcribes audio into text and then the transcribed text is fed into a text NER model.
4.1 Evaluation Metrics
We employed the F1 score metric as it is commonly used for spoken NER Shon et al. (2022); Benaicha et al. (2024), which evaluates an unordered list of NE phrases and tag pairs predicted for each sentence. We used 3 toolkits for a more comprehensive comparison, as described in Appendix D.
@@ 4.2 ASR Models
We employed two baseline models fine-tuned for ASR on VietMed published by Le-Duc (2024): an acoustic monolingual pre-trained w2v2-Viet and an acoustic multilingual pre-trained XLSR-53-Viet model. w2v2-Viet model was pre-trained from scratch on 1204h of Vietnamese data. For the XLSR-53-Viet model, continued pre-training on 1204h of Vietnamese starting with XLSR-53 Conneau et al. (2021) was performed. Both have the same number of parameters (118M) and were fine-tuned on the same training set. Their WERs on the test set are 29.0% and 28.8% respectively.
@@ 4.3 NER Models
Table 2 shows the statistics of various pre-trained monolingual and multilingual models we consider to fine-tune on our dataset. To our knowledge, these are the best pre-trained models that achieved state-of-the-art results on various downstream tasks in the Vietnamese language, including NER.
Monolingual encoder models: PhoBERT_base, PhoBERT_large, PhoBERT_base-v2 Nguyen and Tuan Nguyen (2020), ViDeBERTa_base Tran et al. (2023).
Monolingual sequence-to-sequence (seq2seq) models: BARTpho Tran et al. (2022), ViT5 Phan et al. (2022),
Multilingual encoder models: XML-R_base, XML-R_large Conneau et al. (2020).
Multilingual seq2seq models: mBART-50 Tang et al. (2020).
4.3.1 Seq2seq Training for NER Task
Following the approach proposed by Phan et al. (2021) and later adopted by ViT5 Phan et al. (2022), we formulated the sequence tagging task as a sequence-to-sequence task by training the models to generate tags of labels before and after an entity token. In cases where the models fail to follow the mentioned format for an entity token, we use an “exception” tag, which will be later ignored during metric calculation, as the label.
4.3.2 Training Hyperparameters
We used HuggingFace Transformers Wolf et al. (2019) for fine-tuning pre-trained models for the NER task. Vietnamese input sentences can be represented in either syllable or word level as described by Truong et al. (2021). However, we only employed word-level settings to train NER models. All our NER experiments were done by using the default hyperparameters by HuggingFace.
@@ 5 Experimental Results
Table 3: NER results on reference text of test set. The metrics shown are Precision, Recall, and overall micro F1 score. Results by entity types are shown in Tables 5-24 in the Appendix.
NER ASR Prec. Rec. F1
@@
Table 4: NER results on ASR output of test set for different NER and ASR models. Metrics shown are Precision, Recall, and overall micro F1 score. Results by entity types are shown in Tables 25-44 in the Appendix.
Table 3 and 4 show results of NER using various pre-trained models. We observe that there was a performance drop in all models when evaluated on ASR transcripts, as expected due to the noisy nature of ASR output.
1. Pre-trained multilingual models outperformed monolingual models, if multilingual models overcome the capacity dilution: The pre-trained monolingual model PhoBERT_base-v2 outperformed other monolingual models, at 0.74 of F1 score on reference text, and 0.57 on ASR output. Despite having fewer parameters than PhoBERT_large, it performed similarly, likely due to more pre-training data. The pre-trained multilingual model XLM-R_large achieved the best performance with an F1 score of 0.74 on reference text and 0.58 on ASR output, while XLM-R_base performed worse than PhoBERT_base-v2. This gap is explained by the larger pre-training data (2.5TB multilingual data for XLM-R vs. 140GB monolingual data for PhoBERT_base-v2). Our results with PhoBERT_base-v2 and XLM-R_large confirmed that pre-trained multilingual representations improve performance on medical spoken NER tasks, similar to other language-specific downstream tasks by Conneau et al. (2020); Liu et al. (2020). However, multilingual models may face a Transfer-dilution Trade-off Conneau et al. (2020), where they lack the capacity to learn effective multilingual representations. In other words, for a fixed sized model, the per-language performance decreases as we increase the number of languages Gurgurov et al. (2024). To address this trade-off, multilingual models should possess sufficient capacity, necessitating an adequately large model size Chen and Chen (2024). This is evident in the performance comparison between PhoBERT_base-v2 and XLM-R_base, as seen in other language-specific downstream tasks by Conneau et al. (2020); Arivazhagan et al. (2019).
2. Encoder-based models outperform seq2seq models: The best seq2seq model, BARTpho, achieved F1 scores of 0.68 on reference text and 0.53 on ASR output. Encoders generally performed better than seq2seq models, possibly because seq2seq’s generative nature is less suited for classification tasks like NER.
3. Multi-lingual pre-training of the acoustic model does not affect cascaded NER performance
As expected by the similar WERs for the acoustic pre-trained monolingual model w2v2-Viet and the multilingual model XLSR-53-Viet, all NER models show comparable F1 scores, precision, and recall. This indicates that in addition to overall WER the models do not differ significantly in the recognition accuracy of medical NEs. Non-cascaded models might have advantages in utilising the additional pre-training data for the downstream task.
6 Error Analysis
We performed an error analysis using the best-performing models.
Figure 1: Scatter plot of NER results on reference text by entity types using various pre-trained language models and evaluation variants, created by Tables 5-24.
Figure 2: Scatter plot of NER results on ASR output by entity types using various pre-trained language models and evaluation variants, created by Tables 25-44.
6.1 Quantitative
We provide a detailed error analysis for each entity type across best models, utilizing three evaluation toolkits. The results are summarized in Tables 5-24 and Tables 25-44, with corresponding visual representations in the scatter plots shown in Figure 1 and Figure 2.
The top NER models showed high accuracy in recognizing OCCUPATION and TRANSPORTATION entities in both reference text and ASR output. Despite TRANSPORTATION having only 35 total and 15 unique samples, the best models performed well, likely due to the semantic clarity and predictable spans of these entities.
In contrast, PREVENTIVEMED showed higher misrecognition rates, despite sufficient sample size mitigating class imbalance. This may stem from two factors. First, preventive medicine terms often overlap with general medical terminology, making it difficult for the model to distinguish them from DRUGCHEMICAL or TREATMENT concepts. For example, "vaccination" is frequently misclassified as a therapeutic intervention (TREATMENT) in sentences that describe its role in disease prevention (PREVENTIVEMED). Also, models frequently struggle to differentiate between "vaccination" and "vaccine" (DRUGCHEMICAL). Second, preventive medicine involves long-term health strategies often expressed in non-clinical or non-standardized language, complicating entity recognition.
6.2 Qualitative
A common error was confusion between LOCATION and ORGANIZATION, due to the inherent ambiguity where the same entity can function as either depending on context. An organization-related entity may be labelled as LOCATION if it implies a patient visited there, but this inference requires external knowledge about the entity. Another confusion involved DRUGCHEMICAL and FOODDRINK. Both categories share similar names, descriptors, and consumption contexts (e.g., caffeine, alcohol, sugar). Insufficient context length often causes errors, especially with ambiguous terms like "vitamin," "cordyceps," or "sea daffodils," which can refer to both supplements and nutrients depending on context. Another case is DIAGNOSTICS, TREATMENT, and SURGERY. For example, a "biopsy" can be both a diagnostic and treatment, while "radiation therapy" may be linked to surgery.
A common error in NER involves incorrect entity spans, which fall into two types: (1) correct label but wrong span, and (2) wrong label but correct span. The first type often occurs with multi-word entities in the medical domain, like ORGANIZATION, LOCATION, or DISEASESYMPTOM. For example, "high blood pressure" (B-DISEASESYMPTOM, I-DISEASESYMPTOM, I-DISEASESYMPTOM) may be misrecognized as "high blood" (B-DISEASESYMPTOM, I-DISEASESYMPTOM, O), keeping the meaning but shortening the span. The second type occurs when compound-word entities are split, such as "vagina cells" (B-ORGAN, I-ORGAN) being misrecognized as "vagina" and "cells" (B-ORGAN, B-ORGAN).